6) This question is part of Exercise 3-4 on page 62 of CLRS, but the letters arent all the same. Answer TRUE or FALSE for each of the following statements. TRUE means that the statement is TRUE for any asymptotically positive functions f(n) and g(n). Otherwise, answer FALSE. You dont have to Prove or Disprove these statements... but you should learn how to do that. f(n)- O( g(n)) implies g(n) O( f(n)) b) This is part a in Exercise 3-4. This is part b in Exercise 3-4. This is part g in Exercise 3-4. This is part h in Exercise 3-4. a) fin) + g(n)-Θ( min(f(n), g(n)) ) f(n) = Θ( f(n/2) ) c) d) fin) + 0( f(n) )= Θ( f(n) ) Part d of Question 6 warrants some explanation. What it asks is whether f(n) plus any function that is o(Nn)) must always be Θ( f(n) ), for any asymptotically positive function f(n).7) Prove that for anyb> 1, n = (o( bn ) The Handout CommonFunctions.pdf, which is posted on Piazza under Resources-> General Resources, provides you with an approximation of n! using Stirlings formula that you can use to prove this. That Handout also gives you some information about floor, ceiling and logarithms, that weve already discussed in class. It mentions this problem, but it doesnt prove it.
Answer #1

As per Chegg answering guidelines, I'm answering only the first 4 sub-parts:-

6)

a) FALSE

Take f(n) = 1 and g(n) = n .

Here f(n) = O(g(n)) but g(n) is not O(f(n)).

b) FALSE

As, f(n) + g(n) = θ(max(f(n) g(n)))

c) FALSE

Let f(n) = 22n then f(n/2) = 2n and 22n is not O(2n).

d) TRUE